# On the Reasoning Abilities of Masked Diffusion Language Models

### v1

@@ Appendix D Theoretical gadgets
This section contains various theoretical gadgets that are used in the proofs of the main results. Not all of these are novel, and some are modified restatements from their original sources.
@@ D.1 Positional Encodings
In that vein, it is interesting to consider what is the minimal amount of positional information that the model has to be provided with for it to be able to construct the useful positional encodings. This is described by the following lemma. At a high level, it says that our positional encodings require the binary encodings of the relevant numbers, along with any modular and division operations performed on them. Addition, thresholding, and multiplication by a power of two, in contrast, can be performed by the MLPs in the transformer model.
@@ D.2 Useful attention patterns
      Sub-case 2c: ${{{{{\bm{q}}}}_{{n}}^{\top}{{{\bm{k}}}}_{{{n}}^{\prime}}}}\leq\log{2}({{\texttt{p}}}+1)$. In this case, all intermediate computations are representable in ${{\mathbb{F}}}_{{\texttt{p}}}$ analogously to the case 2a. The attention score is therefore preserved.
@@
The transformer $\mathcal{T}$}} has to convert the binary representation ${\texttt{B}}\left({{n}}\right)$}} of n{{n}} contained in across $\lceil\log{{N}}\rceil$ positions in the input string into a single $\lceil\log{{N}}\rceil$-dimensional binary vector in the residual stream. This is done as follows:
@@ D.3 Neural network constructions
More, concretely, define

### v2

@@ Appendix D Theoretical gadgets
This section contains various theoretical gadgets that are used in the proofs of the main results. Not all of these are novel; some are modified restatements from their original sources.
@@ D.1 Positional Encodings
In that vein, it is interesting to consider what the minimal amount of positional information is that the model has to be provided with for it to be able to construct the useful positional encodings. This is described by the following lemma. At a high level, it says that our positional encodings require the binary encodings of the relevant numbers, along with any modular and division operations performed on them. Addition, thresholding, and multiplication by a power of two, in contrast, can be performed by the MLPs in the transformer model.
@@ D.2 Useful attention patterns
      Sub-case 2c: ${{{{{\bm{q}}}}_{{n}}^{\top}{{{\bm{k}}}}_{{{n}}^{\prime}}}}\leq-\log{2}({{\texttt{p}}}+1)$. In this case, all intermediate computations are representable in ${{\mathbb{F}}}_{{\texttt{p}}}$ analogously to the case 2a. The attention score is therefore preserved.
@@
The transformer $\mathcal{T}$}} has to convert the binary representation ${\texttt{B}}\left({{n}}\right)$}} of n{{n}} spread across $\lceil\log{{N}}\rceil$ positions in the input string into a single $\lceil\log{{N}}\rceil$-dimensional binary vector in the residual stream. This is done as follows:
@@ D.3 Neural network constructions
More concretely, define